Grade 11Mechanics

A body of mass 0.5 kg is placedon the ground.The coefficient of friction between the mass and the ground is 0.2. A horizontal pulling force F=(2sint) is applied on the mass. Find the velocity of the body whne its acceleration becomes zero forthe first time after start.

Amber Agrawal
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
net force = F-f = 2sin t – f
a=0
=>(2sint – f )/0.5=0
=>sint = ½
=>t= π/6
 
a=4sint-0.2
v=∫adt
=-4cost-0.2t| (from 0 to π/6)
=-2√3-0.2π/6+4
=0.43 m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)